OpenStudy (anonymous):

J’Wynn rides his bike 21over4 miles each day. What equation can be used to find r, the number of miles he will ride in 8 days?

OpenStudy (thesecret20111):

The answer would be r=(21/4)*8

OpenStudy (thesecret20111):

Because the total number of miles he would travel in 8 days would be 21/4 multiplied by 8

OpenStudy (anonymous):

so what is the answer ?

OpenStudy (thesecret20111):

Try and work it out.. 21 divided by 4 then that answer multiplied by 8 will equal r

OpenStudy (anonymous):

42?

OpenStudy (thesecret20111):

Yes, very good! But wait the answer is actually asking the equation for working this out.. So put a s the answer r=(21/4)*8

OpenStudy (anonymous):

that's not one of the ansers

OpenStudy (thesecret20111):

Ok is there one that says r=21/4*8?

OpenStudy (anonymous):

no none of them start with r and it is 2 and 1/4

OpenStudy (anonymous):

2 hole and 1/4 one the side

OpenStudy (thesecret20111):

Post all the options and I might be able to understand the question.

OpenStudy (anonymous):

21over4 ÷ 8 = r 21over4 ÷ r = 8 8 × 21over4 = r 8 × r = 21over4

OpenStudy (thesecret20111):

Ok go with this on 8 × 21over4 = r This is the same as r=8 x 21over4 as it doesn't matter what side the = is on, for example if p=6+1 the answer would be p=7 however if if was 6+1=p the answer would still be p=7.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i am going to ask a other one ok

OpenStudy (thesecret20111):

Ok